How to solve equations with two variables on both sides?
E.g.
2n = x
Q: How to make "n" by it's self
Would x be 1/2x
or x/2
or 2/x

Answer by jim_thompson5910(35256)

2n = x

Divide both sides by 2 to get...

2n/2 = x/2

n = x/2


So the answer is n = x/2